OpenStudy (anonymous):

Heather's family bought tickets at a local carnival. Her brother took one-fifth the number of tickets that her cousin took, which was half the number Heather took. If her brother took ten tickets, how many tickets did Heather take?

OpenStudy (anonymous):

@AwesomeAries can u please help

OpenStudy (anonymous):

5.

OpenStudy (anonymous):

All you have to do is read

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

@Rain_Dashie

OpenStudy (anonymous):

This are the choices it gives ten twenty-five fifty one hundred

OpenStudy (texaschic101):

her brother took 1/5, which is half of what heather took. Her brother took 10 1/5 of what is 10 1/5x = 10 x = 10/(1/5) x = 10 * 5/1 x = 50 Her brother took 50 and Heather took half, so Heather took 25 ??

OpenStudy (texaschic101):

you might better get a second opinion...this question is a little bit confusing

OpenStudy (anonymous):

thanks the both of you so much!!!

OpenStudy (texaschic101):

I think I got this wrong....her brother took 1/5, which is half of what heather took. 1/5 of what is 10 1/5x = 10 x = 10/(1/5) x = 10 * 5 x = 50 so if 50 is half of what Heather took, then Heather took 100.

OpenStudy (texaschic101):

you WILL need a second opinion on this..my brain hurts from thinking too much..lol

OpenStudy (anonymous):

yeah i understand LOL

OpenStudy (texaschic101):

wait...I think I got this... Her brother took 1/5 of what her cousin took. And her brother got 10. and 1/5 of 50 is 10, so her cousin got 50 tickets. This is half the number Heather took, so Heather took 50 x 2 = 100 tickets.

OpenStudy (anonymous):

Wow. Much problem. Such difficult. Many wow.

OpenStudy (anonymous):

YEP!!